# Byzantine-resilient Federated Learning With Adaptivity to Data Heterogeneity

### v3

@@ Appendix C Proof of Lemma 1
$\displaystyle\leqslant\frac{8C_{\alpha}}{(2C_{\alpha}-1)^{2}}\sum_{\bm{z}_{i}\in\mathcal{Z}\setminus\mathcal{Z}^{{}^{\prime}}}\alpha_{i}\mathbb{E}\left\lVert\bm{z}_{i}\right\rVert^{2}+\frac{2\epsilon^{2}}{\left(2C_{\alpha}-1\right)^{2}}$ (37)
@@
Please refer to Appendix E. ∎
@@
$\displaystyle\quad+\frac{8C_{\alpha}}{(2C_{\alpha}-1)^{2}}\sum_{m\in\mathcal{M}\setminus\mathcal{B}}\alpha_{m}\mathbb{E}\left\lVert\bm{z}_{m}^{t}-\nabla F(\bm{w}^{t})\right\rVert^{2}$ (38)
@@
$\displaystyle=\frac{1}{K^{t}}\sum_{k=1}^{K^{t}}\mathbb{E}\Big{\{}\Big{\lVert}\nabla F_{m}(\bm{w}_{m}^{t,k-1};\xi_{m}^{t,k})-\nabla F_{m}(\bm{w}_{m}^{t,k-1})+\nabla F_{m}(\bm{w}_{m}^{t,k-1})-\nabla F(\bm{w}^{t})\big{\rVert}^{2}\Big{\}}$ (39d)
$\displaystyle=\frac{1}{K^{t}}\sum_{k=1}^{K^{t}}\mathbb{E}\Big{\{}\left\lVert\nabla F_{m}(\bm{w}_{m}^{t,k-1};\xi_{m}^{t,k})-\nabla F_{m}(\bm{w}_{m}^{t,k-1})\right\rVert^{2}+\left\lVert\nabla F_{m}(\bm{w}_{m}^{t,k-1})-\nabla F(\bm{w}^{t})\right\rVert^{2}$
$\displaystyle\quad+2\big{\langle}\nabla F_{m}(\bm{w}_{m}^{t,k-1};\xi_{m}^{t,k})-\nabla F_{m}(\bm{w}_{m}^{t,k-1}),\nabla F_{m}(\bm{w}_{m}^{t,k-1})-\nabla F(\bm{w}^{t})\big{\rangle}\Big{\}}$ (39e)
where the inequality (39c) holds because of the Cauchy-Schwarz inequality $(a_{1}+a_{2}+a_{3}+\dots+a_{n})^{2}\leqslant n(a_{1}^{2}+a_{2}^{2}+a_{3}^{2}+\dots+a_{n}^{2})$. With Assumptions 3 and 4, there are $\mathbb{E}\left\{\nabla F_{m}(\bm{w}_{m}^{t,k-1};\xi_{m}^{t,k})-\nabla F_{m}(\bm{w}_{m}^{t,k-1})\right\}=0$ and $\mathbb{E}\left\{\left\lVert\nabla F_{m}(\bm{w}_{m}^{t,k-1};\xi_{m}^{t,k})-\nabla F_{m}(\bm{w}_{m}^{t,k-1})\right\rVert^{2}\right\}\leqslant\sigma^{2}$, then (39) leads to
@@
$\displaystyle=\frac{1}{K^{t}}\sum_{k=1}^{K^{t}}\mathbb{E}\Big{\{}\big{\lVert}\nabla F_{m}(\bm{w}_{m}^{t,k-1})-\nabla F(\bm{w}_{m}^{t,k-1})+\nabla F(\bm{w}_{m}^{t,k-1})-\nabla F(\bm{w}^{t})\big{\rVert}^{2}\Big{\}}+\sigma^{2}$ (40b)
$\displaystyle\leqslant\frac{2}{K^{t}}\sum_{k=1}^{K^{t}}\mathbb{E}\left\{\left\lVert\nabla F(\bm{w}_{m}^{t,k-1})-\nabla F(\bm{w}^{t})\right\rVert^{2}\right\}+\frac{2}{K^{t}}\sum_{k=1}^{K^{t}}\mathbb{E}\left\{\left\lVert\nabla F_{m}(\bm{w}_{m}^{t,k})-\nabla F(\bm{w}_{m}^{t,k})\right\rVert^{2}\right\}+\sigma^{2}$ (40c)
@@
To further upper bound the right-hand side of (40), since $\theta_{m}^{t,k}=\left\lVert\nabla F_{m}(\bm{w}_{m}^{t,k})-\nabla F(\bm{w}_{m}^{t,k})\right\rVert^{2}$, we only need to inspect the bound of $\sum_{k=1}^{K^{t}}\mathbb{E}\left\{\left\lVert\nabla F(\bm{w}_{m}^{t,k-1})-\nabla F(\bm{w}^{t})\right\rVert^{2}\right\}$, which is given as follows
@@
  • the inequality (41d) can be derived from Assumptions 3, 4 and 5.
Combining (20), (C), (40) and (41), we obtain the following inequality
$\displaystyle\mathbb{E}\left\{\left\lVert\bm{z}^{t}-\nabla F(\bm{w}^{t})\right\rVert^{2}\right\}\leqslant\Delta^{t}+\frac{8\sigma^{2}(C_{\alpha})^{2}}{(2C_{\alpha}-1)^{2}}+\frac{2\epsilon^{2}}{\left(2C_{\alpha}-1\right)^{2}}$ (42)
@@
Appendix D Proof of Lemma 2
@@
$\displaystyle\mathbb{E}\left\{\left\lVert\bm{z}^{t}\right\rVert^{2}\right\}\leqslant\frac{8C_{\alpha}}{(2C_{\alpha}-1)^{2}}\sum_{m\in\mathcal{M}\setminus\mathcal{B}}\alpha_{m}\mathbb{E}\left\lVert\bm{z}_{m}^{t}\right\rVert^{2}+\frac{2\epsilon^{2}}{\left(2C_{\alpha}-1\right)^{2}}$ (43)
For the term $\mathbb{E}\left\{\left\lVert\bm{z}_{m}^{t}\right\rVert^{2}\right\}$, $m\in\mathcal{M}\setminus\mathcal{B}$, according to Assumptions 3, 4 and 5, we have
@@
$\displaystyle\leqslant G^{2}+\sigma^{2}$ (44c)
@@
$\displaystyle\mathbb{E}\left\{\left\lVert\bm{z}^{t}\right\rVert^{2}\right\}\leqslant\frac{8(C_{\alpha})^{2}(G^{2}+\sigma^{2})}{(2C_{\alpha}-1)^{2}}+\frac{2\epsilon^{2}}{\left(2C_{\alpha}-1\right)^{2}}$ (45)
@@
$\displaystyle 0\leqslant\mathbb{E}\left\{\left\lVert\bm{z}^{t}\right\rVert^{2}\right\}\leqslant\frac{8(C_{\alpha})^{2}(G^{2}+\sigma^{2})}{(2C_{\alpha}-1)^{2}}+\frac{2\epsilon^{2}}{\left(2C_{\alpha}-1\right)^{2}}$ (46)
@@
Appendix E Proof of Lemma 3
@@
$\displaystyle\sum_{i=1}^{M}\frac{S_{i}}{\sum_{j=1}^{M}S_{j}}\left\lVert\bm{z}_{\epsilon}^{*}-\bm{z}_{i}\right\rVert\leqslant\sum_{i=1}^{M}\frac{S_{i}}{\sum_{j=1}^{M}S_{j}}\left\lVert\bm{z}^{*}-\bm{z}_{i}\right\rVert+\epsilon$ (47)
@@
$\displaystyle\sum_{i=1}^{M}\alpha_{i}\left\lVert\bm{z}_{\epsilon}^{*}-\bm{z}_{i}\right\rVert\leqslant\sum_{i=1}^{M}\alpha_{i}\left\lVert\bm{z}^{*}-\bm{z}_{i}\right\rVert+\epsilon$ (48)
@@
$\displaystyle\sum_{i=1}^{M}\alpha_{i}\left\lVert\bm{z}^{*}-\bm{z}_{i}\right\rVert=\inf_{\bm{y}}\sum_{i=1}^{M}\alpha_{i}\left\lVert\bm{y}-\bm{z}_{i}\right\rVert\leqslant\sum_{i=1}^{M}\alpha_{i}\left\lVert\bm{z}_{i}\right\rVert$ (49)
@@
$\displaystyle\sum_{\bm{z}_{i}\in\mathcal{Z}^{{}^{\prime}}}\alpha_{i}\left\lVert\bm{z}_{\epsilon}^{*}-\bm{z}_{i}\right\rVert\geqslant\sum_{\bm{z}_{i}\in\mathcal{Z}^{{}^{\prime}}}\alpha_{i}\left(\left\lVert\bm{z}_{i}\right\rVert-\left\lVert\bm{z}_{\epsilon}^{*}\right\rVert\right)$ (50)
$\displaystyle\sum_{\bm{z}_{i}\in\mathcal{Z}\setminus\mathcal{Z}^{{}^{\prime}}}\alpha_{i}\left\lVert\bm{z}_{\epsilon}^{*}-\bm{z}_{i}\right\rVert\geqslant\sum_{\bm{z}_{i}\in\mathcal{Z}\setminus\mathcal{Z}^{{}^{\prime}}}\alpha_{i}\left(\left\lVert\bm{z}_{\epsilon}^{*}\right\rVert-\left\lVert\bm{z}_{i}\right\rVert\right)$ (51)
@@
$\displaystyle\quad-2\sum_{\bm{z}_{i}\in\mathcal{Z}\setminus\mathcal{Z}^{{}^{\prime}}}\alpha_{i}\left\lVert\bm{z}_{i}\right\rVert$ (52)
@@
$\displaystyle\left\lVert\bm{z}_{\epsilon}^{*}\right\rVert\leqslant\frac{2\sum_{\bm{z}_{i}\in\mathcal{Z}\setminus\mathcal{Z}^{{}^{\prime}}}\alpha_{i}\left\lVert\bm{z}_{i}\right\rVert+\epsilon}{1-2\sum_{\bm{z}_{i}\in\mathcal{Z}^{{}^{\prime}}}\alpha_{i}}$ (53)
@@
$\displaystyle=\frac{8\left(\sum_{\bm{z}_{i}\in\mathcal{Z}\setminus\mathcal{Z}^{{}^{\prime}}}\alpha_{i}\right)^{2}}{\left(1-2\sum_{\bm{z}_{i}\in\mathcal{Z}^{{}^{\prime}}}\alpha_{i}\right)^{2}}\left(\sum_{\bm{z}_{i}\in\mathcal{Z}\setminus\mathcal{Z}^{{}^{\prime}}}\frac{\alpha_{i}}{\sum_{\bm{z}_{j}\in\mathcal{Z}\setminus\mathcal{Z}^{{}^{\prime}}}\alpha_{j}}\left\lVert\bm{z}_{i}\right\rVert\right)^{2}+\frac{2\epsilon^{2}}{\left(1-2\sum_{\bm{z}_{i}\in\mathcal{Z}^{{}^{\prime}}}\alpha_{i}\right)^{2}}$ (54c)
$\displaystyle\leqslant\frac{8\left(\sum_{\bm{z}_{i}\in\mathcal{Z}\setminus\mathcal{Z}^{{}^{\prime}}}\alpha_{i}\right)\sum_{\bm{z}_{i}\in\mathcal{Z}\setminus\mathcal{Z}^{{}^{\prime}}}\alpha_{i}\left\lVert\bm{z}_{i}\right\rVert^{2}+2\epsilon^{2}}{\left(1-2\sum_{\bm{z}_{i}\in\mathcal{Z}^{{}^{\prime}}}\alpha_{i}\right)^{2}}$ (54d)
@@
$\displaystyle\left\lVert\bm{z}_{\epsilon}^{*}\right\rVert^{2}\leqslant\frac{8C_{\alpha}}{\left(2C_{\alpha}-1\right)^{2}}\sum_{\bm{z}_{i}\in\mathcal{Z}\setminus\mathcal{Z}^{{}^{\prime}}}\alpha_{i}\left\lVert\bm{z}_{i}\right\rVert^{2}+\frac{2\epsilon^{2}}{\left(2C_{\alpha}-1\right)^{2}}$ (55)

### v4

@@ Appendix D Proof of Lemma 1
$\displaystyle\leqslant\frac{8C_{\alpha}}{(2C_{\alpha}-1)^{2}}\sum_{\bm{z}_{i}\in\mathcal{Z}\setminus\mathcal{Z}^{{}^{\prime}}}\alpha_{i}\mathbb{E}\left\lVert\bm{z}_{i}\right\rVert^{2}+\frac{2\epsilon^{2}}{\left(2C_{\alpha}-1\right)^{2}}.$ (52)
@@
Please refer to Appendix F. ∎
@@
$\displaystyle\quad+\frac{8C_{\alpha}}{(2C_{\alpha}-1)^{2}}\sum_{m\in\mathcal{M}\setminus\mathcal{B}}\alpha_{m}\mathbb{E}\left\lVert\bm{z}_{m}^{t}-\nabla F(\bm{w}^{t})\right\rVert^{2},$ (53)
@@
$\displaystyle=\frac{1}{K^{t}}\sum_{k=1}^{K^{t}}\mathbb{E}\Big\{\Big\lVert\nabla F_{m}(\bm{w}_{m}^{t,k-1};\xi_{m}^{t,k})-\nabla F_{m}(\bm{w}_{m}^{t,k-1})$
$\displaystyle\qquad\qquad\qquad\quad+\nabla F_{m}(\bm{w}_{m}^{t,k-1})-\nabla F(\bm{w}^{t})\big\rVert^{2}\Big\}$ (54d)
$\displaystyle=\frac{1}{K^{t}}\sum_{k=1}^{K^{t}}\mathbb{E}\Big\{\left\lVert\nabla F_{m}(\bm{w}_{m}^{t,k-1};\xi_{m}^{t,k})-\nabla F_{m}(\bm{w}_{m}^{t,k-1})\right\rVert^{2}$
$\displaystyle\quad+\left\lVert\nabla F_{m}(\bm{w}_{m}^{t,k-1})-\nabla F(\bm{w}^{t})\right\rVert^{2}$
$\displaystyle\quad+2\big\langle\nabla F_{m}(\bm{w}_{m}^{t,k-1};\xi_{m}^{t,k})-\nabla F_{m}(\bm{w}_{m}^{t,k-1}),$
$\displaystyle\qquad\qquad\qquad\qquad\quad\nabla F_{m}(\bm{w}_{m}^{t,k-1})-\nabla F(\bm{w}^{t})\big\rangle\Big\},$ (54e)
where the inequality (54c) holds because of the Cauchy-Schwarz inequality $(a_{1}+a_{2}+a_{3}+\dots+a_{n})^{2}\leqslant n(a_{1}^{2}+a_{2}^{2}+a_{3}^{2}+\dots+a_{n}^{2})$. With Assumptions 9 and 10, there are $\mathbb{E}\left\{\nabla F_{m}(\bm{w}_{m}^{t,k-1};\xi_{m}^{t,k})-\nabla F_{m}(\bm{w}_{m}^{t,k-1})\right\}=0$ and $\mathbb{E}\left\{\left\lVert\nabla F_{m}(\bm{w}_{m}^{t,k-1};\xi_{m}^{t,k})-\nabla F_{m}(\bm{w}_{m}^{t,k-1})\right\rVert^{2}\right\}\leqslant\sigma^{2}$, then (54) leads to
@@
$\displaystyle=\frac{1}{K^{t}}\sum_{k=1}^{K^{t}}\mathbb{E}\Big\{\big\lVert\nabla F_{m}(\bm{w}_{m}^{t,k-1})-\nabla F(\bm{w}_{m}^{t,k-1})$
$\displaystyle\qquad\qquad\qquad+\nabla F(\bm{w}_{m}^{t,k-1})-\nabla F(\bm{w}^{t})\big\rVert^{2}\Big\}+\sigma^{2}$ (55b)
$\displaystyle\leqslant\frac{2}{K^{t}}\sum_{k=1}^{K^{t}}\mathbb{E}\left\{\left\lVert\nabla F(\bm{w}_{m}^{t,k-1})-\nabla F(\bm{w}^{t})\right\rVert^{2}\right\}$
$\displaystyle\quad+\frac{2}{K^{t}}\sum_{k=1}^{K^{t}}\mathbb{E}\left\{\left\lVert\nabla F_{m}(\bm{w}_{m}^{t,k})-\nabla F(\bm{w}_{m}^{t,k})\right\rVert^{2}\right\}+\sigma^{2},$ (55c)
@@
To further upper bound the right-hand side of (55), since $(\theta_{m}^{t,k})^{2}=\left\lVert\nabla F_{m}(\bm{w}_{m}^{t,k})-\nabla F(\bm{w}_{m}^{t,k})\right\rVert^{2}\leqslant\theta^{2}$, we only need to inspect the bound of $\sum_{k=1}^{K^{t}}\mathbb{E}\left\{\left\lVert\nabla F(\bm{w}_{m}^{t,k-1})-\nabla F(\bm{w}^{t})\right\rVert^{2}\right\}$, which is given as follows
@@
  • the inequality (56d) can be derived from Assumptions 9, 10 and 12.
Combining (3), (D), (55) and (56), we obtain the following inequality
$\displaystyle\mathbb{E}\left\{\left\lVert\bm{z}^{t}-\nabla F(\bm{w}^{t})\right\rVert^{2}\right\}\leqslant\Delta^{t}+\frac{8\sigma^{2}(C_{\alpha})^{2}}{(2C_{\alpha}-1)^{2}}+\frac{2\epsilon^{2}}{\left(2C_{\alpha}-1\right)^{2}}.$ (57)
@@
Appendix E Proof of Lemma 2
@@
$\displaystyle\mathbb{E}\left\{\left\lVert\bm{z}^{t}\right\rVert^{2}\right\}\leqslant\frac{8C_{\alpha}}{(2C_{\alpha}-1)^{2}}\sum_{m\in\mathcal{M}\setminus\mathcal{B}}\alpha_{m}\mathbb{E}\left\lVert\bm{z}_{m}^{t}\right\rVert^{2}+\frac{2\epsilon^{2}}{\left(2C_{\alpha}-1\right)^{2}}.$ (58)
For the term $\mathbb{E}\left\{\left\lVert\bm{z}_{m}^{t}\right\rVert^{2}\right\}$, $m\in\mathcal{M}\setminus\mathcal{B}$, according to Assumptions 9, 10 and 12, we have
@@
$\displaystyle\leqslant G^{2}+\sigma^{2},$ (59c)
@@
$\displaystyle\mathbb{E}\left\{\left\lVert\bm{z}^{t}\right\rVert^{2}\right\}\leqslant\frac{8(C_{\alpha})^{2}(G^{2}+\sigma^{2})}{(2C_{\alpha}-1)^{2}}+\frac{2\epsilon^{2}}{\left(2C_{\alpha}-1\right)^{2}}.$ (60)
@@
$\displaystyle 0\leqslant\mathbb{E}\left\{\left\lVert\bm{z}^{t}\right\rVert^{2}\right\}\leqslant\frac{8(C_{\alpha})^{2}(G^{2}+\sigma^{2})}{(2C_{\alpha}-1)^{2}}+\frac{2\epsilon^{2}}{\left(2C_{\alpha}-1\right)^{2}}.$ (61)
@@
Appendix F Proof of Lemma 3
@@
$\displaystyle\sum_{i=1}^{M}\frac{S_{i}}{\sum_{j=1}^{M}S_{j}}\left\lVert\bm{z}_{\epsilon}^{*}-\bm{z}_{i}\right\rVert\leqslant\sum_{i=1}^{M}\frac{S_{i}}{\sum_{j=1}^{M}S_{j}}\left\lVert\bm{z}^{*}-\bm{z}_{i}\right\rVert+\epsilon,$ (62)
@@
$\displaystyle\sum_{i=1}^{M}\alpha_{i}\left\lVert\bm{z}_{\epsilon}^{*}-\bm{z}_{i}\right\rVert\leqslant\sum_{i=1}^{M}\alpha_{i}\left\lVert\bm{z}^{*}-\bm{z}_{i}\right\rVert+\epsilon,$ (63)
@@
$\displaystyle\sum_{i=1}^{M}\alpha_{i}\left\lVert\bm{z}^{*}-\bm{z}_{i}\right\rVert=\inf_{\bm{y}}\sum_{i=1}^{M}\alpha_{i}\left\lVert\bm{y}-\bm{z}_{i}\right\rVert\leqslant\sum_{i=1}^{M}\alpha_{i}\left\lVert\bm{z}_{i}\right\rVert.$ (64)
@@
$\displaystyle\sum_{\bm{z}_{i}\in\mathcal{Z}^{{}^{\prime}}}\alpha_{i}\left\lVert\bm{z}_{\epsilon}^{*}-\bm{z}_{i}\right\rVert\geqslant\sum_{\bm{z}_{i}\in\mathcal{Z}^{{}^{\prime}}}\alpha_{i}\left(\left\lVert\bm{z}_{i}\right\rVert-\left\lVert\bm{z}_{\epsilon}^{*}\right\rVert\right),$ (65)
$\displaystyle\sum_{\bm{z}_{i}\in\mathcal{Z}\setminus\mathcal{Z}^{{}^{\prime}}}\alpha_{i}\left\lVert\bm{z}_{\epsilon}^{*}-\bm{z}_{i}\right\rVert\geqslant\sum_{\bm{z}_{i}\in\mathcal{Z}\setminus\mathcal{Z}^{{}^{\prime}}}\alpha_{i}\left(\left\lVert\bm{z}_{\epsilon}^{*}\right\rVert-\left\lVert\bm{z}_{i}\right\rVert\right).$ (66)
@@
$\displaystyle\quad-2\sum_{\bm{z}_{i}\in\mathcal{Z}\setminus\mathcal{Z}^{{}^{\prime}}}\alpha_{i}\left\lVert\bm{z}_{i}\right\rVert.$ (67)
@@
$\displaystyle\left\lVert\bm{z}_{\epsilon}^{*}\right\rVert\leqslant\frac{2\sum_{\bm{z}_{i}\in\mathcal{Z}\setminus\mathcal{Z}^{{}^{\prime}}}\alpha_{i}\left\lVert\bm{z}_{i}\right\rVert+\epsilon}{1-2\sum_{\bm{z}_{i}\in\mathcal{Z}^{{}^{\prime}}}\alpha_{i}}.$ (68)
@@
$\displaystyle=\frac{8\left(\sum_{\bm{z}_{i}\in\mathcal{Z}\setminus\mathcal{Z}^{{}^{\prime}}}\alpha_{i}\right)^{2}}{\left(1-2\sum_{\bm{z}_{i}\in\mathcal{Z}^{{}^{\prime}}}\alpha_{i}\right)^{2}}\left(\sum_{\bm{z}_{i}\in\mathcal{Z}\setminus\mathcal{Z}^{{}^{\prime}}}\frac{\alpha_{i}}{\sum_{\bm{z}_{j}\in\mathcal{Z}\setminus\mathcal{Z}^{{}^{\prime}}}\alpha_{j}}\left\lVert\bm{z}_{i}\right\rVert\right)^{2}$
$\displaystyle\quad+\frac{2\epsilon^{2}}{\left(1-2\sum_{\bm{z}_{i}\in\mathcal{Z}^{{}^{\prime}}}\alpha_{i}\right)^{2}}$ (69c)
$\displaystyle\leqslant\frac{8\left(\sum_{\bm{z}_{i}\in\mathcal{Z}\setminus\mathcal{Z}^{{}^{\prime}}}\alpha_{i}\right)\sum_{\bm{z}_{i}\in\mathcal{Z}\setminus\mathcal{Z}^{{}^{\prime}}}\alpha_{i}\left\lVert\bm{z}_{i}\right\rVert^{2}+2\epsilon^{2}}{\left(1-2\sum_{\bm{z}_{i}\in\mathcal{Z}^{{}^{\prime}}}\alpha_{i}\right)^{2}},$ (69d)
@@
$\displaystyle\left\lVert\bm{z}_{\epsilon}^{*}\right\rVert^{2}\leqslant\frac{8C_{\alpha}}{\left(2C_{\alpha}-1\right)^{2}}\sum_{\bm{z}_{i}\in\mathcal{Z}\setminus\mathcal{Z}^{{}^{\prime}}}\alpha_{i}\left\lVert\bm{z}_{i}\right\rVert^{2}+\frac{2\epsilon^{2}}{\left(2C_{\alpha}-1\right)^{2}}.$ (70)